OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. (-1,7) and (-10,2) an ordered pair??

OpenStudy (anonymous):

\[ x=\frac{x_1+x_2}2\\ y=\frac{y_1+y_2}2\\ \]

OpenStudy (anonymous):

Can you apply this formula and solve your problem?

OpenStudy (anonymous):

(-1,7) and (-10,2) \[ x=\frac{-10-1}2=-\frac {11}2\\ y=\frac{2+7}2=\frac 9 2\\ \]